OpenStudy (jenniferjuice):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned?

OpenStudy (jenniferjuice):

44 40 76 64

OpenStudy (anonymous):

if \(x\) is the number of men then the number of women is \(100-x\) (since the total is 100)

OpenStudy (anonymous):

the second part tells you \[2x-3(100-x)=20\] and you can solve for \(x\)

OpenStudy (anonymous):

start with \[2x-300+3x=20\] so \[5x-300-20\] making \[5x=320\] divide by \(5\) to find the number of men

OpenStudy (jenniferjuice):

64